# The Malaysian Election Corpus (MECo): Federal and State-Level Election Results from 1955 to 2025

### v2

@@
Thevesh
@@
Thevesh Thevananthan
@@
Amidst this paucity of data, global initiatives such as the Constituency-Level Elections Archive (CLEA)[clea] provide immensely valuable cross-country coverage, including for Malaysian elections. However, they generally focus on federal contests, and within that scope, only on the number of votes won by each candidate (thus omitting information such as the electorate size, voter turnout, unreturned ballots, rejected ballots, etc). Similarly, international turnout or election integrity datasets[icoma2023turnout, norris2014integrity] capture only high-level national indicators. Locally, a number of news and civil society organisations such as Tindak Malaysia[tindak_data], malaysiakini (undi.info), and Sinar Project compile election data to varying degrees of completeness and quality, but these efforts – while laudable for their public service, and valuable as a stopgap measure – typically lack proper data hygiene and standardisation, and are not subjected to systematic validation and review, thus limiting their usefulness for rigorous empirical research and long-term preservation.
In this paper, I address this gap by providing the Malaysian Election Corpus (MECo),[meco_dataset] the first comprehensive open database of Malaysian election results at the federal and state level. MECo is intended as a living resource which provides the go-to empirical foundation for research and journalism on Malaysian elections. The database covers all general elections since the pre-independence general election in 1955, and all 53 by-elections since the 12th federal general election in 2008. In total, it records 14,000 unique candidates representing 110 political parties in 9,704 unique electoral contests from 1955 to 2025.
@@ Methods
Parties and Coalitions. Before discussing the generation of party and coalition UIDs, it is critical to distinguish how ‘party’ is defined in this dataset relative to the EC’s official records. When a candidate puts their name forth for nomination, they are mandated to declare themselves as either representing an entity registered with the EC, or as an independent candidate. However, the EC’s list of registered parties does not distinguish between a single political party and a coalition of parties. For example, as of 30th October 2025, the list contains - without distinction - both the United Malays National Organisation (UMNO) as well as the Barisan Nasional (BN). The former is a ‘party’ in the conventional sense, while the latter is actually a coalition of several parties, including UMNO itself.
@@
    Loss of insight into party dynamics. Political parties which are part of stable coalitions often choose to contest under the coalitional identity. For example, UMNO - Malaysia’s oldest political party - has never contested an election under its own banner, instead contesting under the Alliance flag until 1973, and the BN flag thereafter. This renders it impossible to analyse the political trajectory of UMNO relative to other component parties (especially MIC and MCA) using official data.
@@
For this reason, I made the methodological choice to capture 3 separate variables (see Data Records for complete schema). First, party_on_ballot - which faithfully captures the exact listing of the candidate on the official ballot. Second, party - which captures the true party allegiance of the candidate. Third, coalition - which captures the coalitional allegiance of the party. It was extremely difficult to consolidate this information going back to 1955, especially for state-level results which receive far less attention than federal results - however, I was able to eventually fill in all gaps using a combination of existing scholarly work,[welsh2016elections, hai2017electoral, loh2002democracy, nohlen2001elections] media reports (for modern elections), archived news articles (for older elections), and consultation with colleagues specialising in election history. To the best of my knowledge, MECo is now the only publicly-available dataset with complete data allowing users to distinguish between the 3 dimensions of political allegiance captured here.
@@
Candidates. Candidate UIDs enable the accurate tracking of individual candidates across elections. Beyond the obvious use case of distinguishing candidates with the same name (e.g. the dataset contains 8 unique individuals with the exact same name - ‘Ahmad bin Abdullah’), robust candidate UIDs are particularly important in the Malaysian context, where candidates frequently acquire new honorifics, adopt different formatting conventions, or spell their name differently over time. Two illustrative examples demonstrate the need for candidate UIDs. 8-time Parliamentarian Rafidah Aziz appeared on ballots across 8 elections in forms including:
@@
Once a complete and validated set of candidate UIDs was established, I created a candidate lookup mapping each UID to a cleaned name stripped of titles and honorifics. This lookup was then enriched with three demographic attributes: sex, ethnicity, and date of birth (DOB). Sex and ethnicity were assigned via manual inspection of names, which are nearly perfectly indicative of sex and strongly indicative of ethnic group in the Malaysian context; both variables are complete for 100% of candidates. Where possible, these were verified against public records, particularly for recent elections. DOB were obtained through extensive manual searches across thousands of sources - including biographical directories, news archives, parliamentary profiles, and obituary notices - and recorded in ISO format (YYYY-MM-DD). Reliable birth information could not be found for 33.2% of candidates overall, reflecting limited public documentation in earlier decades. Coverage improves markedly over time: 60% of candidates in GE3 (1969) lack DOB information, falling to 16% by GE10 (1999) and under 5% by GE15 (2022), with remaining gaps concentrated among independent candidates with minimal public profiles.
@@
Consequently, I consider constituency continuity best left to researchers. In fact, it arguably merits dedicated scholarly handling, since electoral lineage has never been rigorously documented or studied in the Malaysian context. That having been said, users who wish to augment MECo with other constituency-based datasets – for example, the subnational statistics published by the Department of Statistics Malaysia – should rely on the intrinsic uniqueness of the date-state-constituency combination as a composite key. This avoids imposing any particular lineage model, while ensuring that external data can be merged cleanly and reproducibly.
@@ 5. $\bullet$
This validation procedure also revealed 115 contests (excluding elections before 1960 which did not report $I$, and the 2006 Sarawak state general election, which did not report $I$, $U$, or $R$) for which the data was digitised accurately based on my source material, but which nevertheless failed to satisfy Equation (1). These errors presumably occurred due to mistakes in data entry which were not caught and corrected during the original publication process. In order to ensure a clean dataset, I applied a standard correction – for all 115 contests failing validation, I adjusted the number of ballots issued (as documented in corrections.csv) such that Equation (1) holds. I chose ballots_issued as the variable to adjust for two reasons. First, until the 1989 amendment of the Elections (Conduct of Elections) Regulations 1981 (implemented at the 1989 federal general election), the number of unreturned ballots was not reported (and thus assumed to be 0). Second, because the number of rejected ballots is very small relative to the number of ballots issued, using rejected ballots as the adjusted variable would have resulted in a much larger correction in percentage terms – specifically, the correction would have required a 20% change to the number of rejected ballots on average (with several exceeding 50%), relative to a 0.6% change to the number of ballots issued (with only one above 5%, and none above 20%).
@@
An interesting trend emerges when I plot the error rate arising from these 115 contests (Figure 5). The error rate was only 0.26% in 1964 (just 1 error), but spiked to nearly 8% in 1969, when elections were severely disrupted by the Emergency. By 1978, error rates stabilised below 2%, and have been constant at 0% since 2008. I posit that this reflects general improvements in data management technology over the decades; I do not have a clear explanation for the absence of errors in 1964.
@@
The Malaysian Election Corpus (MECo) is designed to support a wide range of use cases, ranging from rigorous empirical research to rapid data journalism, and even casual civic technology projects. In the academic realm, MECo – as the first database of its kind for Malaysia – serves as a foundational resource for research in electoral studies, political science, and public policy. Researchers can employ the data to answer important questions about the evolution of Malaysia’s electoral system, employing both cross-sectional and longitudinal analysis.
The data can also be interactively explored via ElectionData.MY. In addition to making the data accessible to non-technical users, the site is intended to serve as the primary channel for continuous updates and enhancements. Future improvements to MECo will be reflected on the site as new versions (and the corresponding source code) are archived on Harvard Dataverse[meco_dataset] and Zenodo[meco_codebase], ensuring that users always have easy access to the freshest data.
@@ 1. $\bullet$
    compile.py, which generates and validates all tabular datasets published on Harvard Dataverse, with the exception of the lookup*.csv files, which were manually curated.
@@ 3. $\bullet$
Author Contributions
TT conceived, designed, and carried out all aspects of this study, including data collection, data curation, analysis, and manuscript preparation.
@@
The author declares no competing interests. The author is not affiliated with any political party, or any organisation which actively participates in or aims to influence the outcome of Malaysian elections.
Acknowledgements
This work did not receive any specific funding. I thank Rosmadi Fauzi and Zulkanain Abdul Rahman for their helpful feedback and suggestions on the dataset and manuscript. I also gratefully acknowledge Nimesha Thevananthan for her assistance with data collection, Danesh Prakash Chacko of Tindak Malaysia for providing historical insights and filling data gaps (particularly for the Sabah state election in 1971), and the MyElectionData volunteers for their assistance with completion and validation of candidate UID harmonisation and candidate demographic data. Finally, I thank the EC for granting access to physical records of historical election results via Pusat Sumber SPR.

### v3

@@
Accepted for Publication This version of the article has been peer-reviewed and accepted for publication in Scientific Data. However, this is not the Version of Record and does not reflect post-acceptance improvements, or any corrections. The Version of Record is available at: doi.org/10.1038/s41597-025-06502-7
@@
Thevesh Thevananthan [^1] Malaya University, W.P. Kuala Lumpur, Malaysia
🖂 thevesh.theva@gmail.com
$^{\text{\Letter}}$
@@
Amidst this paucity of data, global initiatives such as the Constituency-Level Elections Archive (CLEA)[clea] provide immensely valuable cross-country coverage, including for Malaysian elections. However, they generally focus on federal contests, and within that scope, only on the number of votes won by each candidate (thus omitting information such as the electorate size, voter turnout, unreturned ballots, rejected ballots, etc). Similarly, international turnout or election integrity datasets[icoma2023turnout, norris2014integrity] capture only high-level national indicators. Locally, a number of news and civil society organisations such as Tindak Malaysia[tindak_data], malaysiakini (undi.info), and Sinar Project compile election data to varying degrees of completeness and quality, but these efforts—while laudable for their public service, and valuable as a stopgap measure—typically lack proper data hygiene and standardisation, and are not subjected to systematic validation and review, thus limiting their usefulness for rigorous empirical research and long-term preservation.
In this paper, I address this gap by providing the Malaysian Election Corpus (MECo),[meco_dataset] the first complete open database of Malaysian election results at the federal and state level. MECo is intended as a living resource which provides the go-to empirical foundation for research and journalism on Malaysian elections. The database covers all general elections and by-elections since the pre-independence general election in 1955. In total, it records 14,000 unique candidates representing 110 political parties in 9,704 unique electoral contests from 1955 to the present.
@@ Methods
Parties and Coalitions. Before discussing the generation of party and coalition UIDs, it is critical to distinguish how ‘party’ is defined in this dataset relative to the EC’s official records. When a candidate puts their name forth for nomination, they are mandated to declare themselves as either representing an entity registered with the EC, or as an independent candidate. However, the EC’s list of registered parties does not distinguish between a single political party and a coalition of parties. For example, as of 30th October 2025, the list contains—without distinction—both the United Malays National Organisation (UMNO) as well as the Barisan Nasional (BN). The former is a ‘party’ in the conventional sense, while the latter is actually a coalition of several parties, including UMNO itself.
@@
    Loss of insight into party dynamics. Political parties which are part of stable coalitions often choose to contest under the coalitional identity. For example, UMNO—Malaysia’s oldest political party—has never contested an election under its own banner, instead contesting under the Alliance flag until 1973, and the BN flag thereafter. This renders it impossible to analyse the political trajectory of UMNO relative to other component parties (especially MIC and MCA) using official data.
@@
For this reason, I made the methodological choice to capture 3 separate variables (see Data Records for complete schema). First, party_on_ballot—which faithfully captures the exact listing of the candidate on the official ballot. Second, party—which captures the true party allegiance of the candidate. Third, coalition—which captures the coalitional allegiance of the party. It was extremely difficult to consolidate this information going back to 1955, especially for state-level results which receive far less attention than federal results. However, I was able to eventually fill in all gaps using a combination of existing scholarly work,[welsh2016elections, hai2017electoral, loh2002democracy, nohlen2001elections] media reports (for modern elections), archived news articles (for older elections), and consultation with colleagues specialising in election history. To the best of my knowledge, MECo is now the only publicly-available dataset with complete data allowing users to distinguish between the 3 dimensions of political allegiance captured here.
@@
Candidates. Candidate UIDs enable the accurate tracking of individual candidates across elections. Beyond the obvious use case of distinguishing candidates with the same name (e.g. the dataset contains 8 unique individuals with the exact same name—‘Ahmad bin Abdullah’), robust candidate UIDs are particularly important in the Malaysian context, where candidates frequently acquire new honorifics, adopt different formatting conventions, or spell their name differently over time. Two illustrative examples demonstrate the need for candidate UIDs. 8-time Parliamentarian Rafidah Aziz appeared on ballots across 8 elections in forms including:
@@
Once a complete and validated set of candidate UIDs was established, I created a candidate lookup mapping each UID to a cleaned name stripped of titles and honorifics. This lookup was then enriched with three demographic attributes: sex, ethnicity, and date of birth (DOB). Sex and ethnicity were assigned via manual inspection of names, which are nearly perfectly indicative of sex and strongly indicative of ethnic group in the Malaysian context; both variables are complete for 100% of candidates. Where possible, these were verified against public records, particularly for recent elections. DOB were obtained through extensive manual searches across thousands of sources—including biographical directories, news archives, parliamentary profiles, and obituary notices—and recorded in ISO format (YYYY-MM-DD). Reliable birth information could not be found for 33.2% of candidates overall, reflecting limited public documentation in earlier decades. Coverage improves markedly over time: 60% of candidates in GE3 (1969) lack DOB information, falling to 16% by GE10 (1999) and under 5% by GE15 (2022), with remaining gaps concentrated among independent candidates with minimal public profiles.
@@
Consequently, I consider constituency continuity best left to researchers. In fact, it arguably merits dedicated scholarly handling, since electoral lineage has never been rigorously documented or studied in the Malaysian context. That having been said, users who wish to augment MECo with other constituency-based datasets—for example, the subnational statistics published by the Department of Statistics Malaysia—should rely on the intrinsic uniqueness of the date-state-constituency combination as a composite key. This avoids imposing any particular lineage model, while ensuring that external data can be merged cleanly and reproducibly.
@@ 5. $\bullet$
This validation procedure also revealed 114 contests for which the data was digitised accurately based on my source material, but which nevertheless failed to satisfy Equation (1). These errors presumably occurred due to mistakes in data entry which were not caught and corrected during the original publication process. In order to ensure a clean dataset, I applied a standard correction—for all 114 contests failing validation, I adjusted the number of ballots issued (as documented in corrections.csv) such that Equation (1) holds. I chose ballots_issued as the variable to adjust for two reasons. First, until the 1981 amendment of the Elections (Conduct of Elections) Regulations (implemented fully from the 1990 federal general election onwards), the number of unreturned ballots was not reported (and thus assumed to be 0). Second, because the number of rejected ballots is very small relative to the number of ballots issued, using rejected ballots as the adjusted variable would have resulted in a much larger correction in percentage terms—specifically, the correction would have required a 20% change to the number of rejected ballots on average (with several exceeding 50%), relative to a 0.6% change to the number of ballots issued (with only one above 5%, and none above 20%).
@@
An interesting trend emerges when I plot the error rate arising from these 114 contests (Figure 5). The error rate was only 0.26% in 1964 (just 1 error), but spiked to nearly 8% in 1969, when elections were severely disrupted by the Emergency. By 1978, error rates stabilised below 2%, and have been constant at 0% since 2008. I posit that this reflects general improvements in data management technology over the decades; I do not have a clear explanation for the absence of errors in 1964.
@@
The Malaysian Election Corpus (MECo) is designed to support a wide range of use cases, ranging from rigorous empirical research to rapid data journalism, and even casual civic technology projects. In the academic realm, MECo —as the first database of its kind for Malaysia—serves as a foundational resource for research in electoral studies, political science, and public policy. Researchers can employ the data to answer important questions about the evolution of Malaysia’s electoral system, employing both cross-sectional and longitudinal analysis.
The data can also be interactively explored via ElectionData.MY. In addition to making the data accessible to non-technical users, the site is intended to serve as the primary channel for continuous updates and enhancements. Future improvements to MECo will be reflected on the site as new versions are archived on Harvard Dataverse[meco_dataset] and Zenodo[meco_codebase], ensuring that users always have the freshest data.
@@ 3. $\bullet$
Data Availability
All datasets described in this paper are published on Harvard Dataverse[meco_dataset] under a CC0 license. For convenience, the exact same datasets are also mirrored in repositories on Zenodo[meco_codebase] and GitHub. Both contain code and raw source files in addition to the datasets; I use GitHub to version-control active development and maintenance, with substantial updates released on Zenodo.
@@ 1. $\bullet$
    compile.py, which generates and validates all tabular datasets, with the exception of the lookup*.csv files, which were manually curated.
@@ 3. $\bullet$
Acknowledgements
I thank Rosmadi Fauzi and Zulkanain Abdul Rahman for their helpful feedback and suggestions on the dataset and manuscript. I also gratefully acknowledge Nimesha Thevananthan for her assistance with data collection, Danesh Prakash Chacko of Tindak Malaysia for providing historical insights and filling data gaps, and the MyElectionData volunteers for their assistance with completion and validation of candidate UID harmonisation and candidate demographic data. Finally, I thank the EC for granting access to physical election reports via Pusat Sumber SPR.
Author Contributions (CRediT statement)
TT: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Resources, Data curation, Writing - original draft, Writing - review and editing, Visualization, Supervision, Project administration.
Funding
This work did not receive any specific funding.
@@ Acknowledgements
The author declares no competing interests. The author is not affiliated with any political party or election-based organisation.